Q3

Insurance Premiums An insurance company has a simplified method for determining the annual premium for a term life insurance policy. A flat annual fee of $ 150 is charged for all policies plus $2.50 for each thousand dollars of the amount of the policy. For example, a $20,000 policy cost $150 for the fixed fee plus $50, which corresponds to the face value of the policy. ( stated in thousands of dollars ) , determine the function which can be used to compute annual premiums. 


1
Expert's answer
2021-07-29T12:59:21-0400

P= annual premium in dollars

x= face value of the policy (thousands)

$150 for a premium insurance

$2.5 for each thousand dollars

"P=150+(\frac{x}{1000} \times 2.5) \\

P=150 + \frac{x}{400} \\

x=20000 \\

P = 150 + \frac{20000}{400} \\

P=150+50=200"